For the reaction N2(g) + 3 H2(g)→2 NH3(g), ∆rH=−45.94kJ/mol

At 298K,∆rG=−32.8kJ/mol

Estimate∆rG of the same reaction at 0C.
Answer #1

ΔG = -32.8 KJ/mol

ΔH = -45.94 KJ/mol

T = 298 K

use:

ΔG = ΔH - T*ΔS

-32.8 = -45.94 - 298.0 *ΔS

ΔS = -0.0441 KJ/mol.K

ΔS = -44.094 J/mol.K

Now we have:

ΔH = -45.94 KJ/mol

ΔS = -44.094 J/mol.K

= -0.04409 KJ/mol.K

T= 0.0 oC

= (0.0+273) K

= 273 K

use:

ΔG = ΔH - T*ΔS

ΔG = -45.94 - 273.0 * -0.0441

ΔG = -33.9023 KJ/mol

Answer: -33.9 KJ/mol